A spherical conductor of radius 8 cm has a uniform surface charge density of 0.1 nC/m^2. Determine the electric field: a) on the surface of the conductor; b) at a distance of 10 cm from the centre of the conductor.


1
Expert's answer
2022-03-01T11:12:11-0500

The electric field: a) on the surface of the conductor

"E=\frac{q}{4\pi\epsilon_0r^2}=\frac{\sigma}{\epsilon_0}"

"E_1=\frac{0.1*10^{-9}}{8.85*10^{-12}}=11\:\rm V/m"

b) at a distance of 10 cm from the centre of the conductor


"E_2=E_1\frac{r_1^2}{r_2^2}=11*\frac{8^2}{10^2}=7.2\:\rm V/m"